# IA-TIGRIS: An Incremental and Adaptive Sampling-Based Planner for Online Informative Path Planning

### v1

@@ V-B Metrics
The primary metric for evaluating the success of our plan is the information gain along the executed path, quantified as the change in total entropy of the belief space from the initial to the final state. To facilitate interpretation and comparison across tests, we report this as the percent reduction in entropy. Initially, the entropy reduction is zero, as no observations have been made. As the robot gathers information, the entropy decreases, reflecting an improved understanding of the environment. This reduction continues to accumulate throughout the simulation as new observations refine the information map.
@@ V-C IA-TIGRIS Ablations
We conduct ablation experiments to evaluate the impact of incremental and adaptive planning, belief space node embeddings, priority and time dependent rewards, and planner parameters.
@@ V-C1 Incremental and Adaptive Planning Evaluation
To quantify the impact of tree recycling and replanning, we conducted 50 tests in the simulation framework described in Section V-A. Without replanning, the average reduction in entropy was 20.4%. Introducing replanning without tree recycling improved this to 23.1%, a 13.2% increase in performance. Even in the absence of environmental disturbances, replanning improves path quality by addressing mismatches between estimated and actual information gain, particularly as the planning horizon shortens due to decreasing budget. Incorporating incremental planning with tree recycling further increased the reduction in entropy to 25.4%, a 24.5% improvement compared to replanning without tree recycling. This highlights IA-TIGRIS’s to refine and improve the previous global plan.
@@
As expected, the benefits of replanning and refinement become more pronounced when planning time is reduced. When the allowed planning time is reduced to 3 seconds, the performance difference between no replanning and replanning with tree recycling increases to 25.8%. This gap would widen further for even shorter planning times, where initial plans tend to be of lower quality. However, as planning time decreases, a larger portion of the time is allocated to recycling previous plans rather than expanding the tree, leaving less time for refinement. This introduces a tradeoff between path quality and responsiveness; more frequent replanning improves adaptability to new information, while longer planning cycles allow for more extensive trajectory refinement.
V-C2 Belief Space Node Embedding Analysis
We evaluated the impact of belief space embedding on the efficiency of our planning. Calculating the information gain for a trajectory can take up a disproportionate amount of the planning time if not implemented well. We found that in our typical testing environments, using our belief space embedding in the tree nodes reduced the time to compute the information gain by around $80$-$95\%$. This increase in speed only resulted in an additional ${\sim}20$ MB of memory usage for the search tree.
Fig. 8 shows the tree size of the first plan vs time for 100 trials on random simulation environments. Due to the reduction in time to add each new node to the planning tree, the tree size is much larger at any point during planning.
@@ V-C2 Belief Space Node Embedding Analysis
Our embedding also has a profound impact on the speed of recycling the search tree between plans. Previously the information gain at each node would be calculated individually when recycling the tree, with the nodes toward the base of the tree getting evaluated many times. Now, the tree can be updated recursively from the root to the leaves without evaluating the information at any node more than once. When tested in our typical environments, this cut down the time to rebuild a tree from ${\sim}180.93~{}\upmu$s per node to ${\sim}15.48~{}\upmu$s per node.
We also evaluate the impact of belief space embedding by using Google Benchmark [42] on an extremely large environment. For these benchmark tests, the number of nodes in the trajectory was $k=7$, the number of grid cells in the search space was $n=39,841,344$, and the number of grid cells updated by the new node was $j=11,984$. The time to evaluate the information gain for a new node decreased from $43$ ms to $1.9$ ms, a $96\%$ improvement. This led to the size of the search trees to on average be more than double in size.
@@ V-C4 Planner Global Parameter Search
We investigate the effects of the extend distance $\Delta$, the near radius $R$ used in Near()($\cdot$)( ⋅ ), and the pruning radius used in Prune()($\cdot$)( ⋅ ) within IA-TIGRIS, as outlined in Section IV-A. To do this, we perform a global parameter search across 25 randomly generated environments. Each environment was 5 km $\times$ 5 km, with a prior belief map consisting of a random number of Gaussian distributions with varying locations, means, and standard deviations. The planning loop was given a $10$-second planning time, and the agent continued to observe the space until budget was expended.
Figure 10: An isosurface plot showing the results of the global parameter search. Each surface shows given parameters that result in similar average percent of entropy reduction over a set of tests.
Using percent reduction in entropy as the performance metric, we visualize the results in an isosurface plot, shown in Fig. 10, where higher percentages indicate greater information gain. The results suggest that, for the sampled environments, planner performance improves as the extend distance and near radius are larger. If one is much smaller than the other, performance greatly reduced. However, the relationship between pruning radius and performance is less clear, though a slight performance increase is observed as the pruning radius grows.
@@ V-C5 Long-Horizon Planning
For the clustered information map on the left, the long horizon path search over more clusters and achieves a higher overall information gain compared to the short horizon path. This is because the short-horizon planner takes myopic actions, consuming the majority of its budget on the first two clusters. By the time it reaches the third cluster, it exhausts its budget and is unable to reach the last cluster. In contrast, the long horizon planner strategically balances its budget across all four clusters, enabling higher information gain.
For the dense information map on the right, the long-horizon planner generates a smoother path and extends farther into the space compared to the short-horizon planner. However, the long horizon planner allocates a smaller portion of its budget on the highest information gain area near the start position. As a result, the short-horizon planner achieves a slightly better total information gain by prioritizing immediate rewards and concentrating its budget on the highest information cluster close to the starting position.
From these two maps, we observe that the long-horizon planner excels in environments with clustered information where reasoning over long horizons is essential to avoid myopic choices and balance the allocation of budget. The short-horizon planner does perform well in environments with densely distributed information, where following local gradients of information gain can lead to effective solutions. These findings suggest that the impact of the planning horizon on system performance can be environment-dependent. In our testing framework and applications, planning over the entire budget typically yields the best results, as our prior information maps often feature clustered regions of high information gain.
@@ V-D1 Monte-Carlo Tree Search
Here $I(X_{\text{node}})$ denotes the estimated information gain from the node, $\alpha$ denotes the normalization factor which is given by $\frac{B}{v_{\text{desired}}}$, $B$ being the maximum planning budget and $v_{\text{desired}}$ being the desired speed of our UAV. $C$ denotes the exploration weight, and $N_{\text{tree}}$ denotes the number of visits to the tree root node while $N_{\text{node}}$ denotes the number of times the present node has been visited.
After selecting a candidate node, if it has been visited before, it is expanded by applying motion primitives to generate child nodes, growing the tree. Unvisited nodes skip this step. Following expansion, either the unvisited candidate node or one of its children is selected for the simulation phase, where the future values of nodes along the path are estimated to update the total potential information gain. This informs the selection policy in subsequent iterations. Once planning time is exhausted, the path with the highest information gain is returned.
@@ V-E1 Monte Carlo Testing
Our simulated testing environment is a $5000\times 5000$ m square with Gaussian-distributed prior information randomly placed throughout the search space. The number of prior clusters was sampled uniformly between $[4,20]$, with standard deviations between $[60,450]$, and maximum value between $[0.05,0.5]$.
The results of $100$ Monte Carlo tests are shown in Fig. 12. IA-TIGRIS clearly outperforms the other methods, achieving nearly a $40\%$ greater reduction in entropy than the next best method. Early in the simulation, the greedy method initially gains information more quickly, as expected, but this does not translate to better long-term performance. Since our method optimizes for total information gain, it generates paths that maximize information collection over the entire budget. MCTS performed slightly worse than the greedy approach.
@@
We also conducted Monte Carlo tests where either the number of prior clusters or their standard deviation was held constant to analyze how variations in the information map affect planner performance. The results, shown in Fig. 13, include two cases: the upper figure fixes the number of priors, while the lower figure fixes their standard deviation. All other agent and simulation parameters remained unchanged.
@@ VI-A1 Hardware System
VI-A2 Autonomy System
Fig. 16 illustrates the functional system diagram for the real world field test on the DJI M600. The user specifies the initial plan request prior to takeoff. The TIGRIS planner makes an initial plan on that plan request and sends a global path to the waypoint manager. The waypoint manager tracks the current waypoint within the plan and sends the next waypoint to the DJI software development kit, which then sends actuation commands to the motors. The position of the drone is used to calculate the distance from the drone to the ground and sends that distance parameter to the sensor model. The sensor model’s true positive and false positive rate is used to calculate the per-cell entropy updates in the search map manager. The search map manager publishes the current information map, and the replanning node sends an updated plan request to the IA-TIGRIS planner every ten seconds.
@@ VI-A3 Experimental Results
The bottom image of Fig. 14 shows the path selected by IA-TIGRIS in the search area. The figure highlights how the planner dynamically adjusts altitudes over time to balance coverage and sensing resolution, maximizing information gain. Higher altitudes allow for broader area coverage, while lower altitudes provide more detailed observations where needed. Additionally, the planner prioritizes revisiting the three regions of higher uncertainty, recognizing the need for repeated observations reduce entropy. This adaptive strategy ensures that uncertain areas receive sufficient attention to improve the belief map. As a result, the entropy of the information map decreases to near zero by the end of the mission, as shown in Fig. 17(b), indicating that the planner has effectively gathered the necessary information. This behavior demonstrates the planner’s ability to optimize sensing actions, balancing altitude selection, revisit frequency, and exploration to maximize mission success.
@@ VI-B Fixed-wing Deployments
Our proposed approach was extensively tested on the fixed-wing AlareTech TL-1 UAV, shown in Fig. 18. The UAV is equipped with an onboard camera pitched at 10 degrees, which introduces a more challenging planning problem due to the non-holonomic motion model and the camera’s field of view. Over more than 20 flight hours and 100 flights running IA-TIGRIS, we validated our approach with the objective to search for objects of interest in a large search space across a variety of test scenarios, including different terrain types, varying environmental conditions, and diverse target distributions. An example mission from these tests is shown in Fig. 19. In this scenario, the planner was given the search bounds and a designated high-priority region. The resulting flight path prioritized revisiting the high-priority area twice, optimizing sensor use and ensuring maximum information gain. This strategy led to the successful detection of the object of interest, with its estimated position marked by the red dot in the figure.
The map on the upper right in Fig. 19 shows the information map after plan execution was complete. Due to the UAV’s limited budget, the upper right and lower left corners of the map are not searched by the agent. The budget is instead utilized to search over the area of higher priority two times. Compared to the paths in Fig. 14, we observe that the paths for the fixed wing are smoother and have a larger turning radius, demonstrating how IA-TIGRIS respects the motion constraints of the vehicle. We can also see the effect of wind on the path execution, where the flown path shown in green deviates from the planned path shown in yellow. This illustrates the importance of online planning in the cases where this deviation is large or would accumulate over the course of a longer mission and cause the expected observed area to be much different than actual observed area.

### v2

@@ V-B Metrics
The primary metric for evaluating the success of our plan is the information gain along the executed path, quantified as the change in total entropy of the belief map from the initial to the final state. To facilitate interpretation and comparison across tests, we report this as the percent reduction in entropy. Initially, the entropy reduction is zero, as no observations have been made. As the robot gathers information, the entropy decreases, reflecting an improved understanding of the environment. This reduction continues to accumulate throughout the simulation as new observations refine the belief map.
@@ V-C IA-TIGRIS Ablations
We conduct ablation experiments to systematically evaluate the contributions of incremental and adaptive planning, belief map node embeddings, priority and time-dependent rewards, and planner parameters.
@@ V-C1 Incremental and Adaptive Planning Evaluation
To quantify the impact of tree recycling and replanning, we conducted 50 tests in the simulation framework described in Section V-A. Without replanning, the average reduction in entropy was $20.4\!\pm\!2.8$% (mean $\pm$ 95% CI), which would be close to results expected from [21]. Introducing replanning without tree recycling improved this to $23.1\!\pm\!3.3$%, a 13.213.2% relative increase in performance. Even in the absence of environmental disturbances, replanning improves path quality by addressing mismatches between estimated and actual information gain, particularly as the planning horizon shortens due to decreasing budget. Incorporating incremental planning with tree recycling further increased the reduction in entropy to $25.4\!\pm\!3.8$%, a 24.5% relative improvement compared to replanning without tree recycling and a statistically significant difference (p<0.05p\!<\!0.05). This highlights IA-TIGRIS’s ability to refine and improve the previous global plan.
As expected, the benefits of replanning and refinement become more pronounced when planning time is reduced. When the allowed planning time is reduced to 3 seconds, the relative performance difference between no replanning and replanning with tree recycling increases to 25.8%. This gap would widen further for even shorter planning times, where initial plans tend to be of lower quality. However, as planning time decreases, a larger portion of the time is allocated to recycling previous plans rather than expanding the tree, leaving less time for refinement. This introduces a tradeoff between path quality and responsiveness; more frequent replanning improves adaptability to new information, while longer planning cycles allow for more extensive trajectory refinement. In our experiments and applications, we typically set a planning time of 10 seconds, but the exact planning frequency depends on scenario complexity and computational resources.
V-C2 Belief Map Node Embedding Analysis
We evaluated the impact of belief map embedding on planning efficiency. Calculating the information gain for a trajectory can take up a disproportionate amount of the planning time if not implemented well. We found that in our typical testing environments, using our belief map embedding in the tree nodes reduced the time to compute the information gain for each node by around $80$-$95\%$. This increase in speed only resulted in an additional ${\sim}20$ MB of memory usage for the search tree.
The cumulative effect of this speedup is demonstrated in Fig. 8, which compares tree size versus computation time for the first planning iteration, with and without embeddings, across 100 trials on random simulation environments. Due to the reduction in time to add each new node to the planning tree, the tree size is much larger at any point during planning.
@@ V-C2 Belief Map Node Embedding Analysis
Our embedding also has a profound impact on the speed of recycling the search tree between plans. Previously the information gain at each node would be calculated individually when recycling the tree, with the nodes toward the base of the tree getting evaluated many times. Now, the tree can be updated recursively from the root to the leaves without evaluating the information at any node more than once. In our typical environments, this optimization reduced the per-node tree rebuild time from $180.93~\upmu\text{s}~(\sigma=14.7~\upmu\text{s})$ to $15.48~\upmu\text{s}~(\sigma=1.9~\upmu\text{s})$.
We also evaluate the impact of belief map embedding by using Google Benchmark [56] on an extremely large environment. Google Benchmark runs each test repeatedly until the results are statistically stable, automatically accounting for noise and warm-up effects. For these benchmark tests, the number of nodes in the trajectory was $k=7$, the number of grid cells in the search space was $n=39,841,344$, and the number of grid cells updated by the new node was $j=11,984$. The time to evaluate the information gain for a new node decreased from $43$ ms to $1.9$ ms, a $96\%$ improvement. This led to the size of the search trees to on average be more than double in size.
Refer to caption(a)(b)
((a))
((b))
Figure 10: Top: Isosurface plot showing global parameter sweep and resulting entropy reduction. Each surface shows given parameters that result in similar average percent of entropy reduction. Bottom: Slices of the data at fixed extend distance (a) and pruning radius (b).
@@ V-C4 Planner Global Parameter Search
We investigate the effects of the extend distance $\Delta$, the near radius $R$ used in Near()($\cdot$), and the pruning radius used in Prune()($\cdot$) within IA-TIGRIS, as outlined in Section IV-A. To do this, we perform a global parameter search across 25 randomly generated environments. Each environment was 5 km $\times$ 5 km, with a prior belief map consisting of a random number of Gaussian distributions with varying locations, means, and standard deviations. The planning loop was given a $10$-second planning time, and the agent continued to observe the space until budget was expended.
Using the average percent reduction in entropy as the performance metric, we visualize the results in an isosurface plot, shown in Fig. 10, where higher percentages indicate greater information gain. To provide further insight, the bottom of the figure includes two 2D slices of the data: one at an extend distance of 1500 m and the other at a pruning radius of 600 m. These slices reveal the structure of the solution space along the individual parameter dimensions. The results suggest that, for the sampled environments, planner performance improves as the extend distance and near radius are larger. If one is much smaller than the other, performance greatly reduced. However, the relationship between pruning radius and performance is less clear, though a slight performance increase is observed as the pruning radius increases.
@@ V-C5 Long-Horizon Planning
For the clustered belief map on the left, the long horizon path search over more clusters and achieves a higher overall information gain compared to the short horizon path. This is because the short-horizon planner takes myopic actions, consuming the majority of its budget on the first two clusters. By the time it reaches the third cluster, it exhausts its budget and is unable to reach the last cluster. In contrast, the long horizon planner strategically balances its budget across all four clusters, enabling higher information gain.
For the dense belief map on the right, the long-horizon planner generates a smoother path and extends farther into the space compared to the short-horizon planner. However, the long horizon planner allocates a smaller portion of its budget on the highest information gain area near the start position. As a result, the short-horizon planner achieves a slightly better total information gain by prioritizing immediate rewards and concentrating its budget on the highest information cluster close to the starting position.
From these two maps, we observe that the long-horizon planner excels in environments with clustered information where reasoning over long horizons is essential to avoid myopic choices and balance the allocation of budget. The short-horizon planner does perform well in environments with densely distributed information, where following local gradients of information gain can lead to effective solutions. These findings suggest that the impact of the planning horizon on system performance can be environment-dependent. In our testing framework and applications, planning over the entire budget typically yields the best results, as our prior belief maps often feature clustered regions of high information gain.
@@ V-D1 Monte-Carlo Tree Search
Here $I(X_{\text{node}})$ denotes the estimated information gain from the node, $\alpha$ denotes the normalization factor which is given by $\frac{B}{v_{\text{desired}}}$, $B$ being the maximum planning budget and $v_{\text{desired}}$ being the desired speed of our UAV. The scalar $\alpha$ serves two complementary purposes: it aligns the units of the exploitation term with the exploration bonus and it embeds mission‐level context—how much “time” the robot effectively has to collect information—directly into the scoring rule. Otherwise the magnitude of $I(X_{\text{node}})$ can easily dominate the square–root term, leading the search to behave almost greedily. Intuitively, $\frac{B}{v_{\text{desired}}}$ is the maximum flight duration available for the plan; normalizing by this horizon rewards nodes that deliver high information density rather than simply high absolute information. When the operator increases the mission budget $B$, $\alpha$ grows and the exploitation term is damped, encouraging the planner to spend some of the extra allowance on exploration. Conversely, a higher desired speed $v_{\text{desired}}$ (which effectively shortens the decision horizon for a fixed $B$) decreases $\alpha$, sharpening the focus on immediate, high-value observations. Although we did not run a dedicated ablation to isolate the impact of normalization, qualitative tests during early integration were instructive: an unnormalized exploitation term consistently overpowered the exploration bonus, while introducing the $1/\alpha$ factor restored a healthy exploration–exploitation balance under every experiment setting we later report. We leave a more systematic study of normalization effects to future work. $C$ denotes the exploration weight, and $N_{\text{tree}}$ denotes the number of visits to the tree root node while $N_{\text{node}}$ denotes the number of times the present node has been visited.
After selecting a candidate node, if it has been visited before, it is expanded by applying motion primitives to generate child nodes, growing the tree. Unvisited nodes skip this step. Following expansion, either the unvisited candidate node or one of its children is selected for the simulation phase, where the future values of nodes along the path are estimated to update the total potential information gain. This informs the selection policy in subsequent iterations. Once planning time expires, the path with the highest information gain is returned.
@@ V-E1 Monte Carlo Testing
Our simulated testing environment is a $5000\times 5000$ m square with Gaussian-distributed prior information randomly placed throughout the search space. The number of prior clusters was sampled uniformly from the interval $[4,20]$, with standard deviations sampled from $60,450]~\text{m}$, and maximum values from $[0.05,0.5]$.
The results of $100$ Monte Carlo tests are shown in Fig. 12. IA-TIGRIS clearly outperforms the other methods, achieving nearly a $40\%$ greater reduction in entropy than the next best method. Early in the simulation, the greedy method initially gains information more quickly, as expected, but this does not translate to better long-term performance. Since our method optimizes for total information gain, it generates paths that maximize information collection over the entire budget. MCTS performed slightly worse than the greedy approach, likely because it must build a tree of feasible sequences, while the greedy approach can select the globally optimal next action even when it may be far away. This advantage is particularly pronounced in sparse environments where MCTS may not explore distant branches within its computational budget.
@@
We also conducted Monte Carlo tests where either the number of prior clusters or their standard deviation was held constant to analyze how variations in the belief map affect planner performance. The results, shown in Fig. 13, include two cases: the upper figure fixes the number of priors, while the lower figure fixes their standard deviation. All other agent and simulation parameters remained unchanged.
@@ V-E2 Budget Analysis
V-F Comparison Against Learning-based Approach
We additionally compare IA-TIGRIS against a learning-based approach modified from CAtNIPP [16]. CAtNIPP learns a policy to choose sampling locations from neighboring nodes in a pre-built graph. The original approach adopts a Gaussian Process-based information representation and defines information gain as the reduction of the traces of the covariance matrices. We modified the algorithm to use our discrete grid-based representation, changed the reward to reflect information gain from the entire trajectory to the graph node, extended the state space from $(x,y)$ to $(x,y,\psi)$, and applied the identical camera parameters as IA-TIGRIS.
We trained the network using a $15000$ m budget on a $5000\times 5000$ m map with randomly generated Gaussian distribution priors. Over $100$ Monte-Carlo tests with a 200-node graph, the average uncertainty reduction was $15.78\!\pm\!1.9$. This performance is $37.8\%$ worse than IA-TIGRIS, as presented in Table I column $3$, and slightly worse than MCTS and greedy. We attribute this poor performance to the sparse graph structure, which makes it difficult for the network to identify neighboring nodes leading to informative regions.
Training with a larger graph of $1600$ nodes deteriorates the results to $9.85\!\pm\!1.2$ despite $48$ hours of training time, likely due to the inability of the network to learn an effective policy with a larger graph. This reveals a fundamental trade-off that while larger graphs provide better spatial coverage, they become increasingly difficult to train effectively. Therefore, the learning-based approach requires a carefully tuned graph size for specific problem settings and extensive training time to learn a good policy, which greatly limits its use cases.
@@ VI-A1 Hardware System
VI-A2 Autonomy System
Fig. 16 illustrates the functional system diagram for the real world field test on the DJI M600. The user specifies the initial plan request prior to takeoff. The IA-TIGRIS planner makes an initial plan on that plan request and sends a global path to the waypoint manager. The waypoint manager tracks the current waypoint within the plan and sends the next waypoint to the DJI software development kit, which then sends actuation commands to the motors. The position of the drone is used to calculate the distance from the drone to the ground and sends that distance parameter to the sensor model. The sensor model’s true positive and false positive rate (derived from the detector testing data) is used to calculate the per-cell entropy updates in the search map manager. The search map manager publishes the current belief map, and the replanning node sends an updated plan request to the IA-TIGRIS planner every ten seconds.
@@ VI-A3 Experimental Results
The bottom image of Fig. 14 shows the path selected by IA-TIGRIS in the search area. The figure highlights how the planner dynamically adjusts altitudes over time to balance coverage and sensing resolution, maximizing information gain. Higher altitudes allow for broader area coverage, while lower altitudes provide more detailed observations where needed. Additionally, the planner prioritizes revisiting the three regions of higher uncertainty, recognizing the need for repeated observations reduce entropy. This adaptive strategy ensures that uncertain areas receive sufficient attention to improve the belief map. As a result, the entropy of the belief map decreases to near zero by the end of the mission, as shown in Fig. 17(b), indicating that the planner has effectively gathered the necessary information. This behavior demonstrates the planner’s ability to optimize sensing actions, balancing altitude selection, revisit frequency, and exploration to maximize mission success.
@@ VI-B Fixed-wing Deployments
Our proposed approach was extensively tested on the fixed-wing AlareTech TL-1 UAV, shown in Fig. 18. The UAV is equipped with an onboard camera pitched at 10 degrees, which introduces a more challenging planning problem due to the non-holonomic motion model and the camera’s field of view. Over more than 20 flight hours and 100 flights running IA-TIGRIS, we validated our approach with the objective to search for objects of interest in a large search space across a variety of test scenarios, including different terrain types, varying environmental conditions, and diverse target distributions. The object detector was trained on both real and synthetic data, while the sensor model was derived purely from synthetic data in order to get a detailed sensor model for a diverse set of range values and environmental conditions.
@@
An example mission from these tests is shown in Fig. 19. In this scenario, the planner was given the search bounds and a designated high-priority region. The resulting flight path prioritized revisiting the high-priority area twice, optimizing sensor use and ensuring maximum information gain. This strategy led to the successful detection of the object of interest, with its estimated position marked by the red dot in the figure.
The map on the upper right in Fig. 19 shows the belief map after plan execution was complete. Due to the UAV’s limited budget, the upper right and lower left corners of the map are not searched by the agent. The budget is instead utilized to search over the area of higher priority two times. Compared to the paths in Fig. 14, we observe that the paths for the fixed wing are smoother and have a larger turning radius, demonstrating how IA-TIGRIS respects the motion constraints of the vehicle. We can also see the effect of wind on the path execution, where the flown path shown in green deviates from the planned path shown in yellow. This illustrates the importance of online planning in the cases where this deviation is large or would accumulate over the course of a longer mission and cause the expected observed area to be much different than actual observed area.